OpenStudy (anonymous):

best answer will be given . If x varies inversely with y and x = 8 when y = 10, find y when x = 6.

OpenStudy (anonymous):

i am having problems solving this.

OpenStudy (anonymous):

I think I just figured this out its real confusing not even sure if i have it right. Inverse variation formula y= k/x substitue 8 for x and 10 for y 10 = k/8 multiply and solve for k 80=k y=80/x substitute 6 for x y=80/6 divide 80 by 6 y=80/13.333 here is a site it doesn't give you the answer but if you plug in the numbers it will tell you if your right or wrong. http://www.mathick.com/mathick.php?topic=inversevariation hope it helps